Bruner Aeronautics has perpetual preferred stock outstanding with a par value of $100. The stock pays a quarterly dividend of $2, and its current price is $80.
a) What is its nominal annual rate of return ?
b) What is its effective annual rate of return ?
Answer #2

. Bruner Aeronautics has perpetual preferred stock outstanding with a par value of $100. The stock pays a quarterly dividend of $2 and its current price is $80.

a. what is norminal annual rate ofreturn? b. What is its effective annual rate of return?

a.         The preferred stock pays $8 annually in dividends. Therefore, its nominal rate of return would be:

Nominal rate of return = $8/$80 = 10%.

Or alternatively, you could determine the security’s periodic return and multiply by 4.

Periodic rate of return = $2/$80 = 2.5%.

Nominal rate of return = 2.5% ´ 4 = 10%.

=====================================

b.   EAR          = (1 + rNOM/4)4 – 1

        = (1 + 0.10/4)4 – 1

        = 0.103813 = 10.3813%.
Answer #1

a. Nominal annual rate of return = Annual Dividend/* current Price
= $2 * 4 / $80 = 10%

b. Effective annual rate of return = (1 + rNOM/4)4 – 1

= (1+ 0.10/4)^4 - 1 = 10.38%